OpenStudy (anonymous):

The x- intercept of a line is -5 and the y- intercept of the line is -2. What is the equation of the line?

OpenStudy (anonymous):

1.y=-5/2x-5 2.y=2/5x+2 3.y=5/2x-2 4.y=-2/5x-2

OpenStudy (anonymous):

@monyvia23

OpenStudy (anonymous):

Are you sure the signs are typed correct?